OpenStudy (anonymous):

Use the discriminant to determine the number and type of solutions the equation has. x2 + 8x + 12 = 0 A. no real solution B. one real solution C. two rational solutions D. two irrational solutions

OpenStudy (anonymous):

D right?

jhonyy9 (jhonyy9):

do you know the discriminant formula ?

OpenStudy (anonymous):

@Mehek14

jhonyy9 (jhonyy9):

D not is right sure

OpenStudy (anonymous):

Well, I know it is either C or D

jhonyy9 (jhonyy9):

what mean rational and irrational ?

OpenStudy (anonymous):

I am not sure

jhonyy9 (jhonyy9):

so than you need to calculi it - how many will be the discriminant ?

OpenStudy (anonymous):

Ummmmmmmmmmm

jhonyy9 (jhonyy9):

@Mehek14 do you see it now this asker Only removal tips on what is the right answer and he or she dont know the way how can solving an exercise sure right

OpenStudy (anonymous):

Can I ask something?

jhonyy9 (jhonyy9):

what is formula of discriminant ?

OpenStudy (anonymous):

I don't know but can I ask something?

jhonyy9 (jhonyy9):

no you have asked this exercise what you need to solve it

OpenStudy (anonymous):

I don't know. Listen, I really don't like how you are handing this. No offense

jhonyy9 (jhonyy9):

how you wan solving this exercise than you dont know formula of discriminant ?

OpenStudy (anonymous):

I was using Mathway.com but idk what the formula is. I know the answer is -2, -6 though but is that irrantial or rational?

jhonyy9 (jhonyy9):

@Abbster2015 i like learn to you how you can solving this exercise but you need learning formula for discriminant first of all - ok. ?

OpenStudy (anonymous):

Fine.

jhonyy9 (jhonyy9):

so than you dont know this formula of discriminant how you wan solve it ? or you dont learned it again ?

OpenStudy (anonymous):

@Qwertty123 Hey jhonny9? I think I am okay. I don't need anymore help.

OpenStudy (anonymous):

Thanks for your help though

jhonyy9 (jhonyy9):

and again what you have learned when the discriminant is greater than zero and is equal zero and is leas than zero how changing the number of roots of a quadratic ?

OpenStudy (anonymous):

Hey, I am okay. But thanks though.

jhonyy9 (jhonyy9):

ok. good luck bye

OpenStudy (anonymous):

Bye :)

OpenStudy (anonymous):

Anyone else willing to help me?

OpenStudy (anonymous):

I know it's either C or D

OpenStudy (anonymous):

I think it is irrational

OpenStudy (qwertty123):

Still need help? x^2 +8x+12=0 a=1,b=8, c=12 Triangle=b^2-4ac=8^2-4*1*12=64-48=?

OpenStudy (anonymous):

Maybe? Here is my question. Is -2, -6 irrational or ratiioal?

OpenStudy (qwertty123):

"The sum of a rational number and an irrational number is irrational." By definition, an irrational number in decimal form goes on forever without repeating (a non-repeating, non-terminating decimal). By definition, a rational number in decimal form either terminates or repeats.

OpenStudy (anonymous):

I have no clue what that means.

OpenStudy (qwertty123):

Look at what the definition says!

OpenStudy (anonymous):

Oh nevermind. It means rational!

OpenStudy (qwertty123):

YEAH! okay so what did you get for the Problem?

OpenStudy (anonymous):

I ended up getting C as the asnwer

OpenStudy (qwertty123):

YEAH! :D

OpenStudy (anonymous):

Thank you so much!

OpenStudy (qwertty123):

You Welcome! :D